OpenStudy (anonymous):

correct these calculations by changing either side of each equation. a. (x+1)-(2x+1) = -x+2

OpenStudy (anonymous):

\[(4p ^{2}-3q)-4(p ^{2}-q)=7q\]

OpenStudy (anonymous):

thats for b.

OpenStudy (anonymous):

c. \[5k(k+1)-(k ^{2}+5)=4k ^{2}\]

OpenStudy (lukebluefive):

a. \[(x+1)-(2x+1) = -x+2\] Remove the parenthesis: \[x+1-2x-1 = -x+2\] Combine like terms: \[x-2x+1-1 = -x+2\]

OpenStudy (lukebluefive):

b. Similarly to part a: \[(4p^2−3q)−4(p^2−q)=7q\] Remove parenthesis: \[4p^2−3q−4p^2+4q=7q\] Combine like terms: \[4p^2−4p^2−3q+4q=7q\]

OpenStudy (lukebluefive):

c. \[5k(k+1)−(k^2+5)=4k^2\] Multiply out: \[5k^2 + 5k−(k^2+5)=4k^2\] Remove parenthesis: \[5k^2 + 5k−k^2-5=4k^2\] Combine like terms: \[5k^2−k^2 + 5k-5=4k^2\]

OpenStudy (lukebluefive):

As you can see, none of these are actually equivalent. You have to change one of the sides (usually a plus or minus or a constant) to make them correct.

OpenStudy (anonymous):

ok :D thanks for helping :D

OpenStudy (lukebluefive):

You're welcome. Glad I could help. ^_^